OpenStudy (shelby1290):

The population of a small town increased 25% two years ago and then decreased by 25% last year. The population is now 4500 people. What was the population before the two changes?

OpenStudy (anonymous):

4800

OpenStudy (shelby1290):

how did you figure that out

OpenStudy (anonymous):

So the population is 4500 now. We know last year the population decreased 25% so we have the population last year - 25% = 4500 so we can write x - 0.25x = 4500 then we solve for x and we get 6000. Then before the population was 6000 we know it increased by 25% so we have x + 0.25x = 6000 so we solve for x and we get 4800

OpenStudy (shelby1290):

how did you find 6000

OpenStudy (shelby1290):

like how did you solve for x

OpenStudy (anonymous):

we have x - 0.25x = 4500 then we can multiply both sides by 4 and we get 4x - x = 18,000 then 3x = 18,000 then divide both sides by 3 and we get x = 6000

OpenStudy (shelby1290):

ah okay

OpenStudy (shelby1290):

and what about from 6000 to 4800

OpenStudy (shelby1290):

how did you solve for x

OpenStudy (anonymous):

x + 0.25x = 6000 then multiply both sides by 4 so 4x + x = 24,000 then 5x=24,000 then divide both sides by 5 to isolate for x and we get x = 4800

OpenStudy (shelby1290):

OHHHHH i see, thank you